When you purchased your house 4 years ago, you took out a $300,000, 30 year FRM loan at 7.5%. You also paid 2 points to get this loan. Now you want to sell your house and buy another one. In order to do so, you must pay off the existing one. The loan came with 1% prepayment penalty. What has been your effective interest rate on this loan? (show work)(use financial calculator)
Answer #1

Solution :

Steps followed to calculate the effective rate on loan ;

Step 1 - Monthly EMI is calculated.

Step 2 - Financial Calculator is been created in excel for 4 years

Step 3 - Total Effective rate is calculated considering 2 points initial loan avail cost and 1 percent prepayment penalty.

Loan Amount $300,000.00 Interest Rate p.a. 7.5 Tenure 30 years C=(r/1-(1+r)^-N)XP C = Monthly Installment, r = monthly rate o

Effective Interest Rate = (Total Cost / 4)*100 = (97,144.57 /4)*100 = 8.10